# Double-Estimation-Friendly Inference for High Dimensional Misspecified Measurement Error Models

### v1

@@ Abstract
In this paper, we introduce an innovative testing procedure for assessing individual hypotheses in high-dimensional linear regression models with measurement errors. This method remains robust even when either the $X$-model or $Y$-model is misspecified. We develop a double robust score function that maintains a zero expectation if one of the models is incorrect, and we construct a corresponding score test. We first show the asymptotic normality of our approach in a low-dimensional setting, and then extend it to the high-dimensional models. Our analysis of high-dimensional settings explores scenarios both with and without the sparsity condition, establishing asymptotic normality and non-trivial power performance under local alternatives. Simulation studies and real data analysis demonstrate the effectiveness of the proposed method.
Keywords: Measurement error; Model misspecification; High dimensional inference; Double Robust Score Function; Clinical trials.
@@ 1 Introduction
Measurement errors are common in the real world. For example, the measurements of indexes of patients in clinical trials that are difficult to measure or that rely on self-reporting are often inaccurate. The validity of any medical study can be potentially affected by measurement error. Moreover, measurement error can also occur in high dimensional modeling scenarios. For instance, consider a case where we have data on a patient’s self-reported weekly alcohol consumption along with the expression levels of tens of genes from the same patient. The self-reported weekly alcohol consumption of the patient can have possible measurement errors. If we want to build a model to infer the relationship between a specific disease and the amount of alcohol consumption, while removing the effect of genes, we need to consider a high dimensional measurement error model. However, most studies on inference of measurement error models assume low dimensional settings with fixed covariate dimensions. As a result, there is still a paucity of methods for inferring measurement error models in high dimensional settings. Li et al. (2021) proposed a corrected decorrelated score test for high dimensional measurement error models based on the linear model assumption.
@@
The inference of high dimensional misspecified models without any measurement errors has been discussed in the literature. One commonly used approach is to construct the score function, that satisfies the double robust property. It means that this type of score function is locally insensitive to the estimation error, and can obtain a consistent estimator with asymptotical normality when either one of the main model or an ansillary model is correctly specified. Chernozhukov et al. (2018) introduced the general framework, and Bühlmann and van de Geer (2015) demonstrated that the de-sparsified Lasso procedure proposed by Van de Geer et al. (2014) can be adapted to the inference of high dimensional misspecified models. However, these works relies on the sparsity conditions on both models, and need stronger conditions for the main models. This work is closely related to the work under the approximate sparsity condition (Belloni et al., 2014). This condition implies that a linear combination of a limited number of variables can approximate the unknown true mean function within a controllable error margin. In contrast, Shah and Bühlmann (2023) introduced a novel property called the ”double-estimation-friendly” (DEF) property, which can be used to test for the significance of a variable $X$ for a response variable $Y$, controlling for other $p$ covariates $\mathbf{Z}$. The DEF property implies that an inferential procedure is valid when either the model for $X$ regressed on $\mathbf{Z}$, or the model for $Y$ regressed on $(X,\mathbf{Z}^{\top})^{\top}$, is correctly specified. The misspecified part does not need the sparsity condition. Shah and Bühlmann (2023) demonstrated that their testing procedure can satisfy the DEF property, even for misspecified linear models and misspecified generalized linear models. The presence of two models is not only common in high-dimensional statistical inference but also has broad real-world applications. For instance, mediation analysis has been applied in various fields such as finance, psychology, communication research, genetics, and epidemiology. In mediation analysis, researchers investigate the mechanism through which an independent variable (predictor) affects a dependent variable (outcome) through a third variable, known as a mediator. This approach involves the introduction of both an outcome model and a mediation model. For further theory and applications, see MacKinnon (2012), Baron and Kenny (1986), Huang et al. (2015), Huang et al. (2014) and Cai and Chen (2022).
Meanwhile, statistical inference of high dimensional models always relies on the requirement of the assumption of sparsity or approximate sparsity, stating that most of the parameters in the model are zero. However, in practice, the sparsity assumption is not checkable and often violated. Zhu and Bradic (2017), Zhu and Bradic (2018a) and Zhu and Bradic (2018b) proposed a valid inferential procedure by utilizing a Dantzing selector type estimator for dense coefficients.
@@
First, we propose a new inference procedure for one component coefficient that satisfies the DEF property, which is valid when either $X$ regressed on $Z$ is a sparse linear model or $Y$ regressed on $(X,Z^{\top})^{\top}$ is a sparse linear model, allowing one of the two models to be misspecified. Specifically, we construct a double robust score function that permits the linear assumption of one of the models to be violated. We provide a theoretical guarantee for the asymptotic normality of the test statistic under mild conditions in both low-dimensional and high-dimensional settings. This method differs from the corrected score method (Ning and Liu, 2017; Bradic et al., 2020), as it also accounts for the measurement error in $X$.
@@
The rest of this paper is organized as follows. Section 2 presents the motivation and the formula of the double robust moment conditions. The correpsonding inference procedures for this type of misspecification are given in Section 2.1 for low dimensional models and Section 2.2 for high dimensional models. Section 3 investigated the inference procedure by relaxing the sparsity condition in the high dimensional setting. We demonstrate the performance of our approach through simulation studies in Section 4 and real data analysis in clinical trials in Section 5. A brief discussion is given in Section 6. Technical lemmas and proofs are presented in the Appendices in Section 7.
@@ 2 Test Statistics Construction
$Y\text{- Model:}\ \ Y_{i}&=X_{i}\beta+g\left(\mathbf{Z}_{i}\right)+\epsilon_{i},\\ X\text{- Model:}\ \ X_{i}&=f(\mathbf{Z}_{i})+\eta_{i}\ \end{aligned}$ (2)
where $\epsilon_{i}\in\mathbb{R}$ is the error term satisfying $\mathbb{E}(\varepsilon_{i})=0$, $Var(\varepsilon_{i})=\sigma^{2}_{\varepsilon}$ and is assumed to be independent of $\mathbf{X}$ and $\mathbf{Z}$. Similarly, $\bm{\eta}$ is random error independent of $\mathbf{U}$ and $\bm{\varepsilon}$. We also assume that $\mathbb{E}(\eta_{i}|\mathbf{Z}_{i})=0$ and $Var(\eta_{i})=\sigma_{\eta}^{2}$.
We are interested in testing
@@
When both (a) and (b) hold, the original parametric null hypothesis (3) can be transformed into the following moment condition:
@@
with $V_{i}:=Y_{i}-W_{i}\beta^{*}$ as a pseudo-response. Denote $\mathbf{V}=[V_{1},\cdots,V_{n}]^{\top}$. One observation is that moment condition (4) are double-robust in the sense that if only one of the models from (a) and (b) holds, the condition is still satisfied. Under the model (a) in which $\mathbf{Z}_{i}^{\top}\bm{\theta}_{0}$ is not the true formula of $f\left(\mathbf{Z}_{i}\right)$, the moment condition can be rewritten as
$\mathbb{E}\{\left(U_{i}+\eta_{i}+f(\mathbf{Z}_{i})-\mathbf{Z}_{i}^{\top}\bm{\theta}_{0}\right)\left(\epsilon_{i}-U_{i}\beta^{*}\right)+\sigma_{U}^{2}\beta^{*}\}=0.$ (5)
@@
$\mathbb{E}\{\left(U_{i}+\eta_{i}\right)\left(\epsilon_{i}-U_{i}+g(\mathbf{Z}_{i})-\mathbf{Z}_{i}^{\top}\bm{\gamma}_{0}\right)+\sigma_{U}^{2}\beta^{*}\}=0.$ (6)
@@
This moment condition is closely related to the corrected score method, which was first proposed by Nakamura (1990), and the score decorrelation method, which was discussed by Ning and Liu (2017). In the case where $\varepsilon_{i}$ is distributed as a Gaussian distribution and the model in equation (2) is corrected, the corrected score equations for $\bm{\gamma}_{0}$ can be defined as:
@@
2.1 Low Dimensional Setting
For now, we have introduced the basic model setting, and the next step is to investigate the construction of empirical test statistic based on the sample. Firstly, we start from the low dimensional setting when $p$ is fixed. To benefit the analysis, we introduce the following notations. Denote $\bm{\Sigma}=E[(X_{1},\mathbf{Z}_{1}^{T})^{T}(X_{1},\mathbf{Z}_{1}^{T})],\bm{\Sigma}_{X,X}=\mathbb{E}(X_{1}^{2}),$ $\bm{\Sigma}_{Z,Z}=E[Z_{1}Z_{1}^{T}],\bm{\Sigma}_{Z,W}=E[Z_{1}W_{1}]$. $\bm{\Sigma}_{Z,W}=\bm{\Sigma}_{Z,X}$. Let $\lambda_{\max}(\bm{\Sigma})$ and $\lambda_{\min}(\bm{\Sigma})$ be the maximal and minimal eigenvalues of $\bm{\Sigma}$.
@@ 2.1 Low Dimensional Setting
In assumption (A1), we assume the relevant random variables are all sub-Gaussian distributed with uniformly upper bounded Orlicz norm. This is commonly used in the literature of asymptotic analysis (Ning and Liu, 2017; Li et al., 2021). It can further to ensure that the terms all have upper bounded fourth-order moment, which is important to check the Lyapunov’s condition.
@@
  • When model (a) holds, we have:
[…]
  • When model (b) holds, we have:
[…]
Here $\sigma^{2}_{X,Z}:=\bm{\Sigma}_{X,X}-\bm{\Sigma}^{\top}_{X,Z}\bm{\Sigma}^{-1}_{Z,Z}\bm{\Sigma}_{X,Z},\sigma^{2}_{1}=E[(\varepsilon_{1}-U_{1}\beta^{*})(f(Z_{1})-Z^{\top}_{1}\bm{\theta}_{0}+\eta_{1}+U_{1})+\sigma^{2}_{U}\beta^{*}]^{2}$ and $\sigma^{2}_{2}=E[\eta_{1}(\varepsilon_{1}-U_{1}\beta^{*}+g(Z_{1})-Z^{\top}_{1}\bm{\gamma}_{0})+\sigma^{2}_{U}\beta^{*}]^{2}$.
When the $f(\mathbf{Z}_{i})=\mathbf{Z}_{i}^{\top}\bm{\gamma}_{0}$ and $g(\mathbf{Z}_{i})=\mathbf{Z}_{i}^{\top}\bm{\theta}_{0}$, we have
@@
The above results imply that our proposed test statistic $T_{DF}$ can detect a local alternative with a rate of $n^{-1/2}$.
2.2 High Dimensional Setting
Next, let’s turn to the high dimensional measurement error model, where $p$ is allowed to be larger than $n$ and grow with $n$.
We still consider test statistic (12). However, the least squared estimator is not valid when $p>n$. Instead, we can make the statistic adaptive for high dimensional settings by switching $\widehat{\bm{\gamma}}$ and $\widehat{\bm{\theta}}$ in (12) to their corresponding penalized estimators. Namely, we consider the penalized linear regressions:
@@ 2.2 High Dimensional Setting
The forms of $T$ and $T_{DF}$ are unchanged. Let $p_{\lambda}$ be one of $p_{\lambda_{\bm{\theta}}}$ and $p_{\lambda_{\bm{\gamma}}}$. $p_{\lambda}$ should satisfy the following conditions:
@@
The above conditions are the same as Assumption 1 in Loh and Wainwright (2015) and are essential for showing the consistency of $\widehat{\bm{\gamma}}$ and $\widehat{\bm{\theta}}$ to $\bm{\gamma}_{0}$ and $\bm{\theta}_{0}$. The Lasso, the SCAD and the MCP penalty all satisfy these conditions. $\lambda_{\bm{\gamma}}$ and $\lambda_{\bm{\theta}}$ are the corresponding tuning parameters for (13) and (14). In Lemma 2, we will check the conditions in Loh and Wainwright (2015), and show the consistency rate of the estimator under the proposed objective function.
@@
This sparsity condition is commonly used in the high dimensional estimation and inference problem, and is crucial to ensure the consistency of $\widehat{\bm{\gamma}}$ and $\widehat{\bm{\theta}}$ (Javanmard and Montanari, 2014; Ning and Liu, 2017; Li et al., 2021). In practice, the sparsity assumption is hard to check and might be violated. We will discuss above methodology that can work when sparsity assumption of one of $\bm{\gamma}_{0}$ and $\bm{\theta}_{0}$ is violated in Section 3.
@@
Assumptions 6 and 7 are used to ensure the variance $\sigma_{X,\mathbf{Z}}^{2}$, $\sigma_{1}^{2}$, and $\sigma_{2}^{2}$ are bounded from below, which guarantees that the test statistic is valid and possesses non-zero efficiency.
@@
Assume condition (A1) - (A2), (C1) - (C7) hold. If $\lambda_{\bm{\gamma}}=D_{1}\log p/n$ and $\lambda_{\bm{\theta}}=D_{2}\log p/n$, for $D_{1}$,$D_{2}$ sufficiently large, then we have that under null hypothesis, $T_{DF}\rightarrow N(0,1)$ if model (a) or (b) holds.
Theorem 3 establishs the asymptotic normality of the proposed method under the high dimensional setting with suitable choice of the tuning parameters. It can also have local power under local alternatives. We have the following results:
@@
  • When model (a) holds, we have:
[…]
  • When model (b) holds, we have:
[…]
Here $\sigma^{2}_{X,Z}:=\bm{\Sigma}_{X,X}-\bm{\Sigma}^{\top}_{X,Z}\bm{\Sigma}^{-1}_{Z,Z}\bm{\Sigma}_{X,Z}$.
Similar to the low dimensional setting, when the models (a) and (b) both hold, $\sigma_{1}=\sigma_{2}$.
@@
By Theorem 4, we could invert test results to construct confidence regions. We first compute $T_{DF}(\beta_{t})$ to be the test statistic under
@@
A $1-\alpha$ confidence region $C_{\alpha}$ is given by $C_{\alpha}:=\left\{\beta_{t}\in\mathbb{R}:\left|T_{DF}(\beta_{t})\right|\leq\Phi^{-1}(1-\alpha/2)\right\},$ where $\Phi(\cdot)$ is the cumulative distribution function of the of a standard normal distribution. Corollary 1 shows the asymptotic validity of the confidence region.
@@
During our simulation and real data analysis, we used grid search to determine the region of $C_{\alpha}$.

### v3

# Double-Estimation-Friendly Inference for High-Dimensional Measurement Error Models with Non-Sparse Adaptability

@@ Abstract
In this paper, we introduce an innovative testing procedure for assessing individual hypotheses in high-dimensional linear regression models with measurement errors. This method remains robust even when either the $X$-model or $Y$-model is misspecified. We develop a double robust score function that maintains a zero expectation if one of the models is incorrect, and we construct a corresponding score test. We show the asymptotic normality of our approach in a low-dimensional setting, and then extend it to the high-dimensional models. Our analysis of high-dimensional settings explores scenarios both with and without the sparsity condition, establishing asymptotic normality and non-trivial power performance under local alternatives. Simulation studies and real data analysis demonstrate the effectiveness of the proposed method.
Keywords: High-dimensional, Statistical inference, Measurement error, Model misspecification, Double-estimation-friendly.
@@ 1 Introduction
Measurement errors are common in the real world. For example, the measurements of indexes of patients in clinical trials that are difficult to measure or that rely on self-reporting are often inaccurate. The validity of any medical study can be potentially affected by measurement error. Moreover, measurement error can also occur in high dimensional modeling scenarios. For instance, consider a case where we have data on a patient’s self-reported weekly alcohol consumption along with the expression levels of tens of genes from the same patient. The self-reported weekly alcohol consumption of the patient can have possible measurement errors. If we want to build a model to infer the relationship between a specific disease and the amount of alcohol consumption, while removing the effect of genes, we need to consider a high dimensional measurement error model. For examples of measurment error models and its application, see Chen et al. (2005), Fuller (2009), Schennach (2016) and Wang et al. (2023). However, most studies on inference of measurement error models assume low dimensional settings with fixed covariate dimensions. As a result, there is still a paucity of methods for inferring measurement error models in high dimensional settings. Li et al. (2021) proposed a corrected decorrelated score test for high dimensional measurement error models based on the linear model assumption.
@@
The inference of high dimensional misspecified models without any measurement errors has been discussed in the literature. One commonly used approach is to construct the score function, that satisfies the double robust property. It means that this type of score function is locally insensitive to the estimation error, and can obtain a consistent estimator with asymptotical normality when either one of the main model or an ansillary model is correctly specified. Chernozhukov et al. (2018) introduced the general framework, and Bühlmann and van de Geer (2015) demonstrated that the de-sparsified Lasso procedure proposed by Van de Geer et al. (2014) can be adapted to the inference of high dimensional misspecified models. However, these works relies on the sparsity conditions on both models, and need stronger conditions for the main models. This work is closely related to the work under the approximate sparsity condition (Belloni et al., 2014). This condition implies that a linear combination of a limited number of variables can approximate the unknown true mean function within a controllable error margin. In practice, the sparsity assumption is not checkable and often violated. In contrast, Zhu and Bradic (2017), Zhu and Bradic (2018a) and Zhu and Bradic (2018b) proposed a valid inferential procedure by utilizing a Dantzing selector type estimator for dense coefficients.
Additionally, Shah and Bühlmann (2023) introduced a novel property called the ”double-estimation-friendly” (DEF) property, which can be used to test for the significance of a variable $X$ for a response variable $Y$, controlling for other $p$ covariates $Z$. The DEF property implies that an inferential procedure is valid when either the model for $X$ regressed on $Z$, or the model for $Y$ regressed on $(X,Z^{\top})^{\top}$, is correctly specified. Shah and Bühlmann (2023) demonstrated that their testing procedure can satisfy the DEF property, even for misspecified linear models and misspecified generalized linear models.
@@
First, we propose a new inference procedure for one component coefficient that satisfies the DEF property, which is valid when either $X$ regressed on $\mathbf{Z}$ is a sparse linear model or $Y$ regressed on $(X,\mathbf{Z}^{\top})^{\top}$ is a sparse linear model, allowing one of the two models to be misspecified. Specifically, we construct a double robust score function that permits the linear assumption of one of the models to be violated. We provide a theoretical guarantee for the asymptotic normality of the test statistic under mild conditions in both low-dimensional and high-dimensional settings. This method differs from the corrected score method (Ning and Liu, 2017; Bradic et al., 2020), as it also accounts for the measurement error in $X$.
@@
The rest of this paper is organized as follows. Section 2 presents the motivation and the formula of the double robust moment conditions. The correpsonding inference procedures for this type of misspecification are given in Section 2.1 for low dimensional models and Section 2.2 for high dimensional models. Section 3 investigated the inference procedure by relaxing the sparsity condition in the high dimensional setting. We demonstrate the performance of our approach through simulation studies in Section 4 and real data analysis in clinical trials in Section 5. A brief discussion is given in Section 6. Technical lemmas and proofs are presented in the Appendix.
@@ 2 Test Statistics Construction
$Y\text{- Model:}\ \ Y_{i}&=X_{i}\beta+g\left(\mathbf{Z}_{i}\right)+\varepsilon_{i},\\ X\text{- Model:}\ \ X_{i}&=f(\mathbf{Z}_{i})+\eta_{i},\ \end{aligned}$ (2)
where $\varepsilon_{i}\in\mathbb{R}$ is the error term satisfying $\mathbb{E}(\varepsilon_{i})=0$, $Var(\varepsilon_{i})=\sigma^{2}_{\varepsilon}$ and is assumed to be independent of $\mathbf{X}$ and $\mathbf{Z}$. Similarly, $\bm{\eta}$ is random error independent of $\mathbf{U}$ and $\bm{\varepsilon}$. We also assume that $\mathbb{E}(\eta_{i}|\mathbf{Z}_{i})=0$ and $Var(\eta_{i})=\sigma_{\eta}^{2}$. We are interested in testing
@@
The presence of two models is not only common in high-dimensional statistical inference but also has broad real-world applications. For instance, mediation analysis has been applied in various fields such as finance, psychology, communication research, genetics, and epidemiology. In mediation analysis, researchers investigate the mechanism through which an independent variable (predictor) affects a dependent variable (outcome) through a third variable, known as a mediator. This approach involves the introduction of both an outcome model and a mediation model. For further theory and applications, see MacKinnon (2012), Baron and Kenny (1986), Huang et al. (2015), Huang et al. (2014) and Cai and Chen (2022).
Our method is motivated by the observation that when both (a) and (b) hold, the original parametric null hypothesis (3) can be transformed into the following moment condition:
@@
with $V_{i}:=Y_{i}-W_{i}\beta^{*}$ as a pseudo-response and $\mathbf{V}=[V_{1},\cdots,V_{n}]^{\top}$. Moment condition (4) are double-robust in the sense that if only one of the models from (a) and (b) holds, the condition is still satisfied. Under the model (a) in which $\mathbf{Z}_{i}^{\top}\bm{\theta}_{0}$ is not the true formula of $f\left(\mathbf{Z}_{i}\right)$, the moment condition can be rewritten as
$\mathbb{E}\{\left(U_{i}+\eta_{i}+f(\mathbf{Z}_{i})-\mathbf{Z}_{i}^{\top}\bm{\theta}_{0}\right)\left(\varepsilon_{i}-U_{i}\beta^{*}\right)+\sigma_{U}^{2}\beta^{*}\}=0.$ (5)
@@
$\mathbb{E}\{\left(U_{i}+\eta_{i}\right)\left(\varepsilon_{i}-U_{i}+g(\mathbf{Z}_{i})-\mathbf{Z}_{i}^{\top}\bm{\gamma}_{0}\right)+\sigma_{U}^{2}\beta^{*}\}=0.$ (6)
@@
This moment condition is closely related to the corrected score method (Nakamura, 1990) and the score decorrelation method (Ning and Liu, 2017). In the case where $\varepsilon_{i}$ is distributed as a Gaussian distribution and the model in equation (2) is corrected, the corrected score equations for $\bm{\gamma}_{0}$ can be defined as:
@@
2.1 Low-Dimensional Setting
Based on (4), we will construct an empirical test statistic based on the sample. Firstly, we start from the low-dimensional setting when $p$ is fixed. To benefit the analysis, we introduce the following notations. Denote $\bm{\Sigma}=E[(X_{1},\mathbf{Z}_{1}^{T})^{T}(X_{1},\mathbf{Z}_{1}^{T})],\bm{\Sigma}_{X,X}=\mathbb{E}(X_{1}^{2}),$ $\bm{\Sigma}_{Z,Z}=E[Z_{1}Z_{1}^{T}],\bm{\Sigma}_{Z,W}=E[Z_{1}W_{1}]$. $\bm{\Sigma}_{Z,W}=\bm{\Sigma}_{Z,X}$. Let $\lambda_{\max}(\bm{\Sigma})$ and $\lambda_{\min}(\bm{\Sigma})$ be the maximal and minimal eigenvalues of $\bm{\Sigma}$.
@@ 2.1 Low-Dimensional Setting
In Condition (A1), we assume the relevant random variables are all sub-Gaussian distributed with uniformly upper bounded Orlicz norm. This is commonly used in the literature of asymptotic analysis (Ning and Liu, 2017; Li et al., 2021). It can further to ensure that the terms all have upper bounded fourth-order moment, which is important to check the Lyapunov’s condition. Condition (A2) is to avoid extreme values of regression functions.
@@
  • When model (a) holds, it follows that $T_{DF}\rightarrow N\left(c\sigma^{-1}_{1}\sigma^{2}_{X,Z},1\right)$ in distribution, where $\sigma^{2}_{X,Z}:=\bm{\Sigma}_{X,X}-\bm{\Sigma}^{\top}_{X,Z}\bm{\Sigma}^{-1}_{Z,Z}\bm{\Sigma}_{X,Z},\sigma^{2}_{1}=E[(\varepsilon_{1}-U_{1}\beta^{*})(f(Z_{1})-Z^{\top}_{1}\bm{\theta}_{0}+\eta_{1}+U_{1})+\sigma^{2}_{U}\beta^{*}]^{2}$.
  • When model (b) holds, it follows that $T_{DF}\rightarrow N\left(c\sigma^{-1}_{2}\sigma^{2}_{X,Z},1\right)$ in distribution, where $\sigma^{2}_{2}=E[\eta_{1}(\varepsilon_{1}-U_{1}\beta^{*}+g(Z_{1})-Z^{\top}_{1}\bm{\gamma}_{0})+\sigma^{2}_{U}\beta^{*}]^{2}$.
The above results imply that our proposed test statistic $T_{DF}$ can detect a local alternative with a rate of $n^{-1/2}$. It is worth noting when $f(\mathbf{Z}_{i})=\mathbf{Z}_{i}^{\top}\bm{\gamma}_{0}$ and $g(\mathbf{Z}_{i})=\mathbf{Z}_{i}^{\top}\bm{\theta}_{0}$, we have
@@
2.2 High-Dimensional Setting
Next, let’s turn to the high-dimensional measurement error model, where $p$ is allowed to be larger than $n$ and grow with $n$.
We still consider test statistic (12). However, the least squared estimator is not valid when $p>n$. Instead, we can make the statistic adaptive for high-dimensional settings by switching $\widehat{\bm{\gamma}}$ and $\widehat{\bm{\theta}}$ in (12) to their corresponding penalized estimators. Namely, we consider the penalized linear regressions:
@@ 2.2 High-Dimensional Setting
The forms of $T$ and $T_{DF}$ are unchanged. Let $p_{\lambda}(\cdot)$ be one of $p_{\lambda_{\bm{\theta}}}(\cdot)$ and $p_{\lambda_{\bm{\gamma}}}(\cdot)$. $p_{\lambda}$ should satisfy the following conditions:
@@
The above conditions are the same as the Assumption 1 in Loh and Wainwright (2015) and are essential for showing the consistency of $\widehat{\bm{\gamma}}$ and $\widehat{\bm{\theta}}$ to $\bm{\gamma}_{0}$ and $\bm{\theta}_{0}$. The Lasso, the SCAD and the MCP penalty all satisfy these conditions. $\lambda_{\bm{\gamma}}(\cdot)$ and $\lambda_{\bm{\theta}}(\cdot)$ are the corresponding tuning parameters for (13) and (14). In Lemma 2, we will check the conditions in Loh and Wainwright (2015), and show the consistency rate of the estimator under the proposed objective function.
@@
This sparsity condition is commonly used in the high-dimensional estimation and inference problem, and is crucial to ensure the consistency of $\widehat{\bm{\gamma}}$ and $\widehat{\bm{\theta}}$ (Javanmard and Montanari, 2014; Ning and Liu, 2017; Li et al., 2021). In practice, the sparsity assumption is hard to check and might be violated. We will propose a methodology that can work when sparsity assumption of one of $\bm{\gamma}_{0}$ and $\bm{\theta}_{0}$ is violated in Section 3.
@@
Conditions (C6) and (C7) are used to ensure the variance $\sigma_{X,\mathbf{Z}}^{2}$, $\sigma_{1}^{2}$, and $\sigma_{2}^{2}$ are bounded from below, which guarantees that the test statistic is valid and possesses non-zero efficiency.
@@
Assume conditions (A1) - (A2), (C1) - (C7) hold. If $\lambda_{\bm{\gamma}}=D_{1}\log p/n$ and $\lambda_{\bm{\theta}}=D_{2}\log p/n$, for $D_{1}$,$D_{2}$ sufficiently large, then it follows that under null hypothesis, $T_{DF}\rightarrow N(0,1)$ in distribution if model (a) or (b) holds.
Theorem 3 establishs the asymptotic normality of the proposed method under the high-dimensional setting with suitable choice of the tuning parameters. It can also have local power under local alternatives. We have the following results:
@@
  • When model (a) holds, it follows that $T_{DF}\rightarrow N\left(c\sigma^{-1}_{1}\sigma^{2}_{X,Z},1\right).$ in distribution, where $\sigma^{2}_{X,Z}:=\bm{\Sigma}_{X,X}-\bm{\Sigma}^{\top}_{X,Z}\bm{\Sigma}^{-1}_{Z,Z}\bm{\Sigma}_{X,Z},\sigma^{2}_{1}=E[(\varepsilon_{1}-U_{1}\beta^{*})(f(Z_{1})-Z^{\top}_{1}\bm{\theta}_{0}+\eta_{1}+U_{1})+\sigma^{2}_{U}\beta^{*}]^{2}$.
  • When model (b) holds, it follows that $T_{DF}\rightarrow N\left(c\sigma^{-1}_{2}\sigma^{2}_{X,Z},1\right)$ in distribution, where $\sigma^{2}_{2}=E[\eta_{1}(\varepsilon_{1}-U_{1}\beta^{*}+g(Z_{1})-Z^{\top}_{1}\bm{\gamma}_{0})+\sigma^{2}_{U}\beta^{*}]^{2}$.
Similar to the low-dimensional setting, when the models (a) and (b) both hold, $\sigma_{1}=\sigma_{2}$.
@@
By Theorem 3 and Theorem 4, we could invert test results to construct confidence regions. We first compute $T_{DF}(\beta_{t})$ to be the test statistic under
@@
The $(1-\alpha)$ confidence region $C_{\alpha}$ is given by
$C_{\alpha}:=\left\{\beta_{t}\in\mathbb{R}:\left|T_{DF}(\beta_{t})\right|\leq\Phi^{-1}(1-\alpha/2)\right\},$
where $\Phi(\cdot)$ is the cumulative distribution function of the of a standard normal distribution. Corollary 1 shows the asymptotic validity of the confidence region. During our simulation and real data analysis, we used grid search to determine the region of $C_{\alpha}$.